Problem 12.23 (Excel Video) Cullumber Company management is considering a project that will require an initial investment of $50,000 and will last for 10 years. No other capital expenditures or increases in working capital are anticipated during the life of the project. What is the annual EBIT that will make the project economically viable if the cost of capital for the project is 10 percent and the firm will depreciate the investment using straight-line depreciation and a salvage value of $0? Assume that the marginal tax rate is 30 percent. (Do not round intermediate calculations. Round final answer to 2 decimal places, e.g. 52.75.) Excel Template (Note: This template includes the problem statement as it appears in your textbook. The problem assigned to you here may have different values. When using this template, copy the problem statement from this screen for easy reference to the values you’ve been given here, and be sure to update any values that may have been pre-entered in the template based on the textbook version of the problem.) EBIT $
Answer #1
NPV = OCF*PVIFA(10,10)-50000
For being economically viable, the NPV should be atleast 0.
Therefore, 0 = [EBIT*(1-t)+Depreciation]*PVIFA(10,10)-50000
Substituting values, we have
0 = (EBIT*0.7+5000)*6.14457-50000
EBIT*0.7*6.14457 = 50000-5000*6.14457
EBIT = (50000-5000*6.14457)/(0.7*6.14457) = $     4,481.81
CHECK:
NPV = (4481.81*0.7+5000)*6.14457-50000 = $             0.01
Almost 0